3.3 In a class of 100 learners, 80 do Mathematics and 35 do Physics.

 If x learners do Mathematics and Physics, and y learners do 

 neither Mathematics nor Physics, find the greatest possible value

 of y.


1
Expert's answer
2021-08-04T11:54:24-0400

Solution:

"n(T)=100,n(M)=80,n(P)=35,
\\n(M\cap P)=x,
\\n(M'\cap P')=y=n(M\cup P)'
\\\Rightarrow n(T)-n(M\cup P)=y
\\\Rightarrow 100-n(M\cup P)=y
\\\Rightarrow n(M\cup P)=100-y"

Using "n(M \cup P)=n(M)+n(P)-n(M\cap P)"

"\Rightarrow 100-y=80+35-x
\\\Rightarrow y=x-15"

Now, "n(M\cap P)\le n(P)"

"\Rightarrow x\le35"

Now for maximum value of "y", "x" must be 35.

So, "y=35-15=20"